OpenStudy (anonymous):

HELP PLEASE ! The length of a rectangle is one unit more than its width. If the area is 30 square units, find the dimensions of the rectangle.

OpenStudy (kkutie7):

area=30 L=w+1

OpenStudy (anonymous):

you said length equals width plus 1?

OpenStudy (kkutie7):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

im so lost

OpenStudy (kkutie7):

oh I for the =0 on the end of the last equation

OpenStudy (kkutie7):

I'm sorry what don't you understand?

OpenStudy (anonymous):

the equation doesnt make sense to me....

OpenStudy (kkutie7):

understand?

OpenStudy (anonymous):

no im sorry im just gonna try some other problems and come back to these later

OpenStudy (kkutie7):

I'll stay here and help you with what you need. You just need to be very specific with what you do or don't understand and I'll explain. Please don't be sorry.

OpenStudy (anonymous):

Okay, thanks but can we work on something else? (a different problem)

OpenStudy (kkutie7):

ok just tag me in it when you post =)

OpenStudy (anonymous):

okay

OpenStudy (kkutie7):

*A=L*W

OpenStudy (kkutie7):

A=(W+1)*W 30=(W+1)*W 30=W^2+W W^2+W-30=0 (x-6)(x+5)=0 x=-6,5 can't have a negative number so we assume 5

OpenStudy (kkutie7):

sorry those x's should be w's

OpenStudy (kkutie7):

L=5+1=6

OpenStudy (kkutie7):

A=30 L=6 W=5 I came back here to fix this. It was brought to my attention I didn't do it right the first time. I apologize.

OpenStudy (anonymous):

its cool thanks